OpenStudy (anonymous):

Solve:|3u - 18| = 0 A.6 B.9 C.{-6, 6} D.{6, 9}

OpenStudy (anonymous):

a

OpenStudy (anonymous):

click good answer when someone answers your questions ffs

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

wrong love its like this |3u-18|=0 3u=18u=9

OpenStudy (anonymous):

Hmm... I keep doin it wrong.

OpenStudy (anonymous):

Yeah, it's A. 3*6=18 18-18= 0 It's A.

OpenStudy (anonymous):

never mind i was the same befire

OpenStudy (anonymous):

I had it right, then did it wrong, now i have it right.

OpenStudy (anonymous):

its not 3u 18u, its 3u - 18 = 0 3u = 18 u = 6

OpenStudy (anonymous):

|3u-18|=0 3u-18=0 u=6

OpenStudy (anonymous):

i knew tht

OpenStudy (anonymous):

Medal? O>O